OpenStudy (turingtest):

Let\[y=x^p(1+x)^q\]find\[y^{(p+q)}\]

OpenStudy (turingtest):

that's the (p+q)th derivative by the way Sorry that I have to go soon, but I look forward to seeing your responses when I return.

OpenStudy (amistre64):

ln(y) = ln(x^p) + ln(1+x)^q y'/y = p/x + q/(1+x) hmm

OpenStudy (anonymous):

i think the answer is just (p+q)!

OpenStudy (anonymous):

since the highest power of x will be x^(p+q), and all the other powers of x will become 0 after P+q derivatives.

OpenStudy (amistre64):

sounds plausible to me

OpenStudy (turingtest):

hint:(Leibniz' theorem/ binomial stuff)

OpenStudy (turingtest):

at least in MIT's solution I'd like to see others

OpenStudy (amistre64):

lol, id like to see mits solution :)

OpenStudy (turingtest):

I'll send it to you....

OpenStudy (accessdenied):

im curious, what class is this problem from?

OpenStudy (turingtest):

OCW calc I

OpenStudy (accessdenied):

Ahh, okay. Thank you. :)